On the first day of the fiscal year, a company issues an $886,000, 7%, 5-year bond that pays semiannual interest of $31,010 ($886,000 x 7% x 1/2), receiving cash of $832,800. Journalize the entry to record the first interest payment and the amortization of the related bond discount using the straight-line method.

If an amount box does not require an entry, leave it blank.
Answer #1

Discount on bonds payable = $886000-832800 = $53200

Account Titles Debit Credit
Interest Expense $       36,330 =31010+5320
        Cash $               31,010 =886000*7%*6/12
        Discount on Bonds Payable $                  5,320 =53200/10
(Being interest expense recorded)